OpenStudy (anonymous):

4√6/√30 please show the steps to solving it

OpenStudy (anonymous):

\[\LARGE \frac{4\sqrt6}{\sqrt{30}}=\frac{4\sqrt6}{\sqrt5 \cdot \sqrt6}\] \[\LARGE \frac{4}{\sqrt5}\cdot \frac{\sqrt5}{\sqrt5}=\frac{4\sqrt5}{5}\] on the second step we rationalized ... because we can't let denominator under the square root :)

OpenStudy (anonymous):

do you understand it?

OpenStudy (anonymous):

Yes Thankyou, can you do this one and do the same thing?

OpenStudy (anonymous):

(2√5+3√7)^2